OpenStudy (anonymous):

A regular hexagon has sides of 2 feet. What is the area of the hexagon?

OpenStudy (anonymous):

Use the formula A = 1/2aP A = 1/2aP I know from experience that you get a 30-60-90 triangle, so just use the ratio to get the apothem. A = 1/2(√3)(12) A = 6√3